find the work done when 1 mole of gas is expanded reversibily isothermal from 5atm to 1 atm at 25degree centigrate
 

Grade:11

1 Answers

Aman
6 years ago
 Wisothermal = 2.303  n R T log P1/P2 
n = no of moles of gas
R = universal gas constant = 8.314 J / mol. K
T = temperature in kelvin (at which isothermal process is occuring) = 25 +273 = 298 K
P1 and P2 = Initial pressure and final pressure. (not necessary to convert into SI units because, its ratio and units will cancel out)
P1 / P2 = 5 
 
Substitute all values, we get    (take log 5 = 0.67)
W = 3988J = 3.9 KJ